OpenStudy (anonymous):

How do I solve this using the multiplication property? 5/2x>-9/4

OpenStudy (lgbasallote):

\[\frac 52 x > -\frac 94\] multiply both sides by \(\large \frac 25\) because it is the reciprocal of \(\large \frac 52\)

OpenStudy (anonymous):

x>-18/20